# Identification and Optimal Nonlinear Control of Turbojet Engine Using Koopman Eigenfunction Model

### v2

@@ 2.1.4 Turbine
Similarly to the compressor, the turbine was represented by its map. Concerning the turbine map, there was no need for a coordinate transformation and the map could be directly interpolated. The turbine corrected gas mass flow $W_{t,corr}$ and the relative temperature difference $\Delta T/T_{t}$ were functions of the turbine corrected spool speed $N_{t,corr}$ and pressure ratio $\Pi_{t}$ generally expressed as
$\left[W_{t,corr},\frac{\Delta T}{T_{t}}\right]=f(N_{t,corr},\Pi_{t}).$ (4)
@@
where $C_{p,g}$ is the heat capacity of gas (1.33 for the combustion products).
@@ 2.1.5 Nozzle
The nozzle model included the pressure losses due to both the friction and contraction. The critical nozzle pressure ratio $\Pi_{n,crit}$ was also accounted for during the exhaust velocity and thrust calculation. The nozzle velocity $v_{n}$, mass flow $W_{n}$, and the thrust $F$ are given as
$v_{n}=\begin{cases}\sqrt{2C_{p,g}\hskip 1.42262ptT_{n}\left(1-\Pi_{n}^{\frac{1-\gamma_{g}}{\gamma_{g}}}\right)\eta_{n}}&\quad\text{for }\Pi_{n}<\Pi_{n,crit}\\ \sqrt{\frac{2\gamma_{g}}{\gamma_{g}+1}RT_{n}}&\quad\text{for }\Pi_{n}\geq\Pi_{n,crit}\end{cases},$ (6)
@@
$F=W_{n}v_{n}-W_{0}MN\sqrt{\gamma_{a}RT_{0}}+p_{n,out}A_{n},$ (8)
where $\eta_{n}$, $T_{n}$ and $p_{n}$ are the adiabatic efficiency, total temperature and total pressure in the nozzle, respectively, $p_{0}$ and $T_{0}$ are the atmospheric pressure and temperature, respectively, $\Pi_{n}=p_{n}/p_{0}$ is the nozzle pressure ratio, $\gamma_{g}$ is the adiabatic exponent of exhaust gases, $R$ is the universal gas constant, $\mu_{n}$ denotes the nozzle flow coefficient, $A_{n}$ is the nozzle outlet area, $W_{0}$ is the engine inlet air mass flow, $MN$ is the Mach number of flight, and $p_{n,out}$ is the exhaust static pressure measured directly aft the nozzle ($p_{n,out}>p_{0}$ for supercritical conditions).
@@ 2.1.6 Shaft dynamics and control system
$\frac{dN}{dt}=\frac{30}{\pi I}\left(\Gamma_{t}\eta_{m}-\Gamma_{c}\right),$ (9)
where the spool speed $N$ is measured in revolution per minute (RPM), $I$ is the shaft polar moment of inertia, and $\eta_{m}$ is the mechanical efficiency, which is usually very close to 1.
@@
Control system consisted of the main spool-speed controller (governor) and fuel flow limiters connected to the min-max protection logic. The fuel flow limits were calculated as functions of the spool speed accounting for the compressor surge and the turbine over-temperature during acceleration, and the compressor choke and combustor blow-out during deceleration. The clamping method was utilized to suppress the integrator wind up effect.
@@ 2.2 Transient solver
For the transient calculation, the constant mass flow (CMF) approach coupled with the Newton-Raphson matching algorithm was utilized. For modeling of single-spool turbojet GTEs, the CMF method is advantageous due to its higher accuracy and suitability for long-term transient simulations, as reported in [49, 50]. For more complex engines, e.g. two-spool turbofans, the method becomes computationally expensive and less accurate due to the unmodeled volume dynamics.
In each time step of the transient computation, the change of the air/gas mass flow induced by the changes of the engine state is assumed to instantaneously propagate through the entire engine. Therefore, in any instance, the mass flow of air/gas entering a particular component must be equal to the mass flow leaving the component. For a single-spool turbojet engine, this matching condition has to be satisfied between the compressor and the turbine, and between the turbine and the nozzle. The goal is to find the values of two state variables of the Newton-Raphson method. These variables are guessed and updated until the convergence, mass flow equality, is reached. These state variables are the compressor $\beta_{c}$ parameter and the turbine pressure ratio $\Pi_{t}$. The matching errors are given by differences between mass flows read off the compressor and the turbine map, and the corresponding mass flows computed based on previous cycle calculations. To update the state variables, the following equation is utilized:
@@
where $e_{m,c}$ and $e_{m,n}$ are the compressor-turbine and turbine-nozzle mass flow errors, respectively, and partial derivatives in the Jacobi matrix were approximated using the small perturbation method and finite differences.
The complete system consists of a controller to be designed, the fuel system which can be accurately represented by a linear model (10), and the GTE which is a nonlinear system that needs to be identified from data. The state variables of the GTE and fuel system are the spool speed and fuel flow, respectively. The input to the fuel system is the fuel-flow command from the controller. As only the engine rotor dynamics need to be treated as nonlinear, the following identification procedures and transformations will be applied only to the spool speed. The fuel system dynamics will later be augmented with the identified GTE model in the Koopman observable subspace.
@@ 3 Koopman operator theory
The Koopman operator theory provides a powerful framework for representation of nonlinear dynamical systems in a new coordinate system, where the dynamics become linear. The operator itself is theoretically infinite-dimensional. In practice however, the finite-rank approximation is of interest and an entire plethora of tools and methods for its discovery was discussed in literature. One way is represented by DMD-based methods described in the introduction. In this paper, the other approach based on a direct identification of Koopman eigenfunctions was considered. Consider the following nonlinear continuous-time dynamical system:
@@
Furthermore, consider a set of nonlinear Lipchitz continuous and Lebesgue integrable observable functions $\mathbf{g}(\textbf{x},u)=\begin{bmatrix}\mathbf{\Psi(x)}&\mathbf{\tilde{u}}(\textbf{x},u)\end{bmatrix}$, $\mathbf{\tilde{u}}(\textbf{x},0)=\textbf{0}$, representing the Koopman lifting functions. Thus, it is assumed that the unforced dynamics can be described using only the observables $\mathbf{\Psi}(\textbf{x})$. The associated Koopman operator $\mathcal{K}^{t}$ is defined by the following composition equation:
@@
Note that there exists an entire family of Koopman operators parameterized by $t$ (sampling period). For the purpose of presenting eigenfunctions, the Koopman generator form is more convenient. When $t\rightarrow 0$, the infinitesimal change of observables is given by a Lie derivative as
@@
Infinitely many Koopman operators $\mathcal{K}^{t}$ can be sampled from a single generator given a time step $t$. For the unforced dynamics, special observables called eigenfunctions are particularly of interest. The Koopman eigenfunction $\varphi(\textbf{x})$ and the corresponding eigenvalue $\lambda$ are defined as [28]
@@
In practice, the theoretical operator $\mathcal{K}$ is approximated by a matrix K, and associated eigen-tuples obtained from data-driven methods such as EDMD. For a system with inputs, this matrix can be further split as $\mathbf{K=\begin{bmatrix}\mathbf{A}&\textbf{B}\end{bmatrix}}$, where A and B are the system and input matrices, respectively. The eigenfunctions and modes can be obtained by diagonalization of A as it represents the unforced (autonomous) dynamics. EDMD is one of the common methods for the data-driven identification of Koopman systems [32]. The continuous-time Koopman linear system is thus given as
@@
In EDMD, a library of nonlinear observable functions is first constructed for lifting of the state dynamics. Subsequently, the finite-rank matrix K is obtained by LS minimization and represents an optimal mapping from one time step to the next across the entire dataset. Finally, diagonalization of the system matrix A is carried out to obtain eigenvalues and eigenfunctions as
@@
However, as these eigenfunctions are constrained to be linear combinations of the same basis $\mathbf{\Psi}$, the limitations are already given by the selection of $\mathbf{\Psi(x)}$. These observables should thus be selected such that the projection $\mathbf{\Phi=V^{-1}\Psi}$ is well-defined for the most prominent eigenfunctions, which is not always guaranteed. Furthermore, EDMD might lead to solutions that are ill-conditioned and possibly contain unstable eigenvalues although the model provides very accurate fit [31]. The selection of optimal EDMD observables is often a trial-and-error process, which can be supported by approximation and learning methods as discussed in section 1.
In the spectral approach, these prominent eigenfunctions are identified directly from data with no need for construction of a global set of observables since the eigenfunctions themselves are the optimal intrinsic coordinates directly connected to the rich underlying dynamics. Moreover, they span an invariant subspace due to their algebraic structure [44, 45]. $\mathbf{\Phi(x)}$ can be expressed as a linear combination of arbitrary basis functions that can adapt to each eigenfunction individually. Thus, there are no constraints given by the projection of eigenfunctions on a common fixed set of observables.
@@
The KPDE is in agreement with the generator formulation (14, 3) and uniquely defines eigenfunctions for a given set of eigenvalues. This is particularly useful for identification of eigenfunctions from data [31].
@@ 4 Proposed approach - SINDy-KEM
Eigenfunctions can be reliably identified for the autonomous dynamics of the system. However, in the case of GTEs, it is rather challenging to measure autonomous nature of the dynamics as mentioned in Section 1.
The idea of direct identification of eigenfunctions from data using regression and KPDE originates in the discussed works of Korda et al. [45] and Kaiser et al. [31]. In this paper, it is further combined with a SINDy algorithm utilized for estimation of the autonomous dynamics. This enables discovery of the underlying dynamics from data gathered during normal engine operation. Thus, the extensive experimental procedures, e.g. linearization around multiple steady-state points for derivation of an LPV model, are bypassed. Furthermore, there is no need for measurements of the component characteristics and the engine geometry, both of which are not always accessible, and simplifying assumptions that might compromise the accuracy when chosen inappropriately.
The workflow includes the identification of the engine governing equation and extraction of the estimated autonomous dynamics, discovery of the optimal set of eigenvalues, corresponding eigenfunctions and eigenmodes, and transform of the nonlinear governing equation to the KEM. Summary is provided in Fig. 1.
@@ 4.1 Identification of the control-affine governing equation
The SINDy model is the most important part of the entire process since it sets the maximum accuracy of the KEM. Thus, it is crucial to properly capture the nonlinear dynamics by SINDy. The dynamics of single-spool GTEs can be generally represented by
@@
where $p_{1t}$ and $T_{1t}$ capture the variations in inlet conditions caused by external effects.
Let us hereafter use the shorter notation $N$ and $W_{f}$ for corrected quantities $N_{corr}$ and $W_{f,corr}$, respectively, given as
$N_{corr}=N\sqrt{\frac{288}{T_{1t}}}\rightarrow N,$ (21)
$W_{f,corr}=W_{f}\sqrt{\frac{288}{T_{1t}}}\frac{101325}{p_{1t}}\rightarrow W_{f},$ (22)
where $p_{1t}$ and $T_{1t}$ are the inlet total pressure and temperature, respectively.
The simplified control-affine model of the GTE is given as
@@
where $f$ and $g$ are the autonomous and input dynamics, respectively, and $N$ will hereafter denote the normalized corrected spool speed for simplicity.
@@
$J_{S}=\left\|\dot{N}-\mathbf{\Xi_{f}}\mathbf{\Theta_{f}}(N,\mathbf{p_{f}})+\mathbf{\Xi_{g}}\mathbf{\Theta_{g}}(N,\mathbf{p_{g}})W_{f}\right\|_{F}^{2}+\alpha\|\mathbf{\Xi}\|_{1},$ (24)
@@
Since also the nonlinear parameters of basis functions are optimized, this problem is solved using gradient descend (GD) methods with adaptive momentum (ADAM) optimizer [51]. Additionally, to promote sparsity of the solution, the L1 regularization is combined with sequential thresholding (ST) [24]. Whenever any weight in $\mathbf{\Xi}$ falls below a specified threshold, it is excluded from the equation. After each ST round, the GD algorithm is initialized again to fine-tune the remaining active terms in the model. This procedure repeats until the suitable trade-off between the complexity, given by the number of active terms, and accuracy is reached.
@@
$\begin{bmatrix}\mathbf{\Xi}\\ \textbf{p}\end{bmatrix}_{k+1}=\begin{bmatrix}\mathbf{\Xi}\\ \textbf{p}\end{bmatrix}_{k}-\frac{\alpha\textbf{v}}{\sqrt{\textbf{s}+\epsilon}},$ (25)
@@
$\textbf{v}=\beta_{1}\textbf{v}+(1-\beta_{1})\nabla J_{S},$ (26)
$\textbf{s}=\beta_{2}\textbf{s}+(1-\beta_{2}){\nabla J_{S}}^{2},$ (27)
$\alpha=\eta_{p}\left(\frac{1-{\beta_{2}}^{i}}{1-{\beta_{1}}^{i}}\right),$ (28)
$\nabla J_{S}=\begin{bmatrix}\frac{\partial J_{S}}{\partial\mathbf{\Xi}}&\frac{\partial J_{S}}{\partial\textbf{p}}\end{bmatrix}^{\textbf{T}},$ (29)
@@
$\nabla J_{0}=K_{0}\frac{|f|}{f}\begin{bmatrix}\frac{\partial f}{\partial\mathbf{\Xi}}&\frac{\partial f}{\partial\textbf{p}}\end{bmatrix}^{\textbf{T}}\bigg{|}_{t=0},$ (31)
@@
The ST is performed according to the following condition:
$\xi(\xi<\text{threshold})=0,\text{for any }\xi\in\mathbf{\Xi}$ (32)
and also according to properties of given basis functions. For instance, if logistic functions (LFs) were utilized, the center shifted too far from the active region would make the function either 0 or 1, which would affectively exclude it or merge it with the other constant functions in the set.
Next step is the extraction of the autonomous dynamics $f(N)$ by setting $W_{f}=0\hskip 2.84526ptkg/s$, and generating a set of autonomous trajectories. The function $f(N)$ is integrated starting from multiple initial conditions until the zero state is reached. This provides an estimate of the true autonomous dynamics of the GTE. It should be noted that the zero state, $N=0$, corresponds to the minimum sustainable physical spool speed, usually in the ground idle (GI) regime, w.r.t. which the normalization of $N$ is performed.

### v6

@@ 2.1.4 Turbine
Similarly to the compressor, the turbine was represented by its map. However, there was no need for a coordinate transformation, and the map could be directly interpolated. The turbine corrected gas mass flow $W_{t,corr}$ and the relative temperature difference $\Delta T/T_{t}$ were functions of the turbine corrected spool speed $N_{t,corr}$ and pressure ratio $\Pi_{t}$ generally expressed as
$\left[W_{t,corr},\frac{\Delta T}{T_{t}}\right]=f_{t}(N_{t,corr},\Pi_{t}).$ (4)
@@
where $W_{t}$ is now the physical turbine mass flow and $C_{p,g}$ is the heat capacity of gas.
@@ 2.1.5 Nozzle
The nozzle model included the pressure losses due to both friction and contraction. The critical nozzle pressure ratio $\Pi_{n,crit}$ was also accounted for during the exhaust velocity and thrust calculation. The nozzle velocity $v_{n}$, mass flow $W_{n}$, and the thrust $F$ are given as
$v_{n}=\begin{cases}\sqrt{2C_{p,g}\hskip 1.42262ptT_{n}\left(1-\Pi_{n}^{\frac{1-\gamma_{g}}{\gamma_{g}}}\right)\eta_{n}}&\text{for }\Pi_{n}<\Pi_{n,crit}\\ \sqrt{\frac{2\gamma_{g}}{\gamma_{g}+1}RT_{n}}&\text{for }\Pi_{n}\geq\Pi_{n,crit}\end{cases},$ (6)
@@
$F=W_{n}v_{n}-W_{0}M_{0}\sqrt{\gamma_{a}RT_{0}}+p_{n,out}A_{n},$ (8)
where $\eta_{n}$, $T_{n}$ and $p_{n}$ are the adiabatic efficiency, total temperature and total pressure in the nozzle, respectively, $p_{0}$ and $T_{0}$ are the atmospheric pressure and temperature, respectively, $\Pi_{n}=p_{n}/p_{0}$ is the nozzle pressure ratio, $\gamma_{g}$ is the adiabatic exponent of exhaust gases, $R$ is the universal gas constant, $\mu_{n}$ denotes the nozzle flow coefficient, $A_{n}$ is the nozzle outlet area, $W_{0}$ is the engine inlet air mass flow, $M_{0}$ is the Mach number of flight, and $p_{n,out}$ is the exhaust static pressure measured directly aft the nozzle ($p_{n,out}>p_{0}$ for supercritical conditions).
@@ 2.1.6 Shaft dynamics and control system
$\frac{\text{d}N}{\text{d}t}=\frac{30}{\pi I}\bigg(\Gamma_{t}\eta_{m}-\Gamma_{c}\bigg),$ (9)
where the spool speed $N$ is measured in revolutions per minute (RPM), $I$ is the shaft polar moment of inertia, and $\eta_{m}$ is the mechanical efficiency, which is usually very close to 1.
@@
The control system consisted of the main spool-speed controller (governor) and fuel flow limiters connected to the min-max protection logic. The fuel flow limits were calculated as functions of the spool speed, accounting for the compressor surge and the turbine over-temperature during acceleration, and the compressor choke and combustor blow-out during deceleration. The clamping method was utilized to suppress the integrator wind-up effect.
@@ 2.2 Transient solver
For the transient calculation, the constant mass flow (CMF) approach coupled with the Newton-Raphson matching algorithm was utilized. For modeling of single-spool turbojet GTEs, the CMF method is advantageous due to its higher accuracy and suitability for long-term transient simulations, as reported in [bib49, bib50]. For more complex engines, e.g., two-spool turbofans, the method becomes computationally expensive and less accurate due to the unmodeled volume dynamics.
In each time step of the transient computation, the change of the air/gas mass flow induced by the changes of the engine state is assumed to instantaneously propagate through the entire engine. Therefore, in any instance, the mass flow of air/gas entering a particular component must be equal to the mass flow leaving the component. For a single-spool turbojet engine, this matching condition has to be satisfied between the compressor and the turbine, and between the turbine and the nozzle. The goal is to find the values of two state variables of the Newton-Raphson method. These variables are guessed and updated until the convergence, mass flow equality, is reached. These state variables are the compressor $\beta_{c}$ parameter and the turbine pressure ratio $\Pi_{t}$. The matching errors are given by differences between mass flows obtained from the compressor and the turbine map, and the corresponding mass flows computed based on previous cycle calculations. To update the state variables, the following equation is utilized:
@@
where $e_{m,c}$ and $e_{m,n}$ are the compressor-turbine and turbine-nozzle mass flow errors, respectively. The partial derivatives in the Jacobi matrix were approximated using the small perturbation method and finite differences.
The complete system is depicted in Fig. 1. It consists of a controller to be designed, the fuel system, which can be accurately represented by the linear model (10), and the GTE, which is a nonlinear system that needs to be identified from data. The state variables of the GTE and fuel system are the spool speed, $N$, and fuel flow, $W_{f}$, respectively. The input to the fuel system is the fuel-flow command from the controller. As only the engine rotor dynamics need to be treated as nonlinear, the following identification procedures and transformations will be applied only to the spool speed dynamics. The linear dynamics of the fuel system will be later augmented with the identified GTE model in the Koopman observable subspace.
Figure 1: Block diagram of the system
@@ 3 Koopman operator theory
The KOT provides a powerful framework for the representation of nonlinear dynamical systems in a new coordinate system, where the dynamics become linear. The operator itself is theoretically infinite-dimensional. In practice, however, the finite-rank approximation is of interest, and an entire plethora of tools and methods for its discovery were discussed in the literature. One category are the DMD-based methods described in the introduction. In this paper, another approach based on a direct identification of Koopman eigenfunctions was considered. Consider the following nonlinear continuous-time dynamical system:
@@
Furthermore, consider a set of nonlinear, Lipschitz continuous, and Lebesgue integrable observable functions:
$\mathbf{g}(\textbf{x},u)$ $=\begin{bmatrix}\mathbf{\Psi(x)}&\mathbf{\tilde{u}}(\textbf{x},u)\end{bmatrix}$,
$\mathbf{\tilde{u}}(\textbf{x},0)$ $=\textbf{0},$
representing the Koopman lifting functions. Thus, it is assumed that the unforced dynamics can be described using only the observables $\mathbf{\Psi}(\textbf{x})$. The associated Koopman operator $\mathcal{K}^{t}$ is defined by the following composition equation:
@@
Note that there exists an entire family of Koopman operators parameterized by $t$ (sampling period). To present eigenfunctions, the Koopman generator form is more convenient. When $t\rightarrow 0$, the infinitesimal change of observables is given by a Lie derivative as
@@
Infinitely many Koopman operators $\mathcal{K}^{t}$ can be sampled from a single generator given a time step $t$. For the unforced dynamics, special observables called eigenfunctions are particularly of interest. The Koopman eigenfunction $\varphi(\textbf{x})$ and the corresponding eigenvalue $\lambda$ are defined as [bib28]
@@
In practice, the infinite-dimensional operator $\mathcal{K}$ is approximated by a matrix K, and associated eigen-tuples obtained from data-driven methods such as EDMD. For a system with inputs, this matrix can be further split as $\mathbf{K=\begin{bmatrix}\mathbf{A}&\textbf{B}\end{bmatrix}}$, where A and B are the system and input matrices, respectively. The eigenfunctions and modes can be obtained by diagonalization of A as it represents the unforced (autonomous) dynamics. EDMD is one of the common methods for the data-driven identification of Koopman systems [bib32]. The continuous-time Koopman linear system is thus given as
@@
In EDMD, a library of nonlinear observable functions is first constructed for the state dynamics lifting. Subsequently, the finite-rank matrix K is obtained by LS minimization and represents an optimal mapping from one time step to the next across the entire dataset. Finally, diagonalization of the system matrix A is carried out to obtain eigenvalues and eigenfunctions as
@@
However, as these eigenfunctions are constrained to linear combinations of the same basis, $\mathbf{\Psi}$, the limitations are already given by the selection of $\mathbf{\Psi(x)}$. These observables should be designed in such a way that the projection $\mathbf{\Phi=V^{-1}\Psi}$ is well-defined for the most prominent eigenfunctions, which is not always guaranteed. Furthermore, EDMD might lead to solutions that are ill-conditioned and possibly contain lightly unstable eigenvalues, although the model itself provides a very accurate fit [bib31]. The selection of optimal EDMD observables is often a trial-and-error process, which can be supported by approximation and learning methods as discussed in Section 1.
In the spectral approach, these prominent eigenfunctions are identified directly from data with no need for the construction of a global set of observables, since the eigenfunctions themselves are the optimal intrinsic coordinates directly connected to the rich underlying dynamics. Moreover, they span an invariant subspace due to their algebraic structure [bib44, bib45]. $\mathbf{\Phi(x)}$ can be expressed as a linear combination of arbitrary basis functions that can adapt to each eigenfunction individually. Thus, there are no constraints given by the projection of eigenfunctions on a common fixed set of observables.
@@
The KPDE agrees with the generator formulation (14, 3) and uniquely defines eigenfunctions for a given set of eigenvalues. This is particularly useful for the identification of eigenfunctions from data [bib31].
@@ 4 Proposed approach - SINDy-KEM
Eigenfunctions can be reliably identified for the autonomous dynamics of the system. However, in the case of GTEs, it is rather challenging to measure the autonomous nature of the dynamics as mentioned in Section 1.
The idea of direct identification of eigenfunctions from data using regression and KPDE originates in the discussed works of Korda et al. [bib45] and Kaiser et al. [bib31]. In this paper, it is further combined with the SINDy algorithm utilized for the estimation of the autonomous dynamics. This enables the discovery of the underlying dynamics from data gathered during normal engine operation. Thus, the extensive experimental procedures, such as linearization around multiple steady-state points for deriving an LPV model, are bypassed. Furthermore, there is no need for measurements of the component characteristics and the engine geometry, both of which are not always accessible, and simplifying assumptions that might compromise the accuracy when chosen inappropriately.
The workflow involves identifying the engine governing equation, extracting the estimated autonomous dynamics, discovering the optimal set of eigenvalues, corresponding eigenfunctions, and eigenmodes, and transforming the nonlinear governing equation into the KEM. Subsequently, the control system is designed and optimized. The summary is provided in Fig. 2.
@@ 4.1 Identification of the control-affine governing equation
The SINDy model is the most important part of the entire process since it sets the maximum accuracy of the KEM. Thus, it is crucial to properly capture the nonlinear dynamics. The dynamics of single-spool GTEs can be generally described as
@@
where $p_{1t}$ and $T_{1t}$ are the inlet total pressure and temperature, respectively, capturing varying flight conditions and external effects.
For a more compact representation, still accounting for varying flight conditions, the corrected quantities are utilized as follows [bib1] (Chapter 8):
$N_{corr}$ $=N_{phys}\sqrt{\frac{288}{T_{1t}}},$ (21)
$W_{f,corr}$ $=W_{f,phys}\sqrt{\frac{288}{T_{1t}}}\frac{101325}{p_{1t}},$ (22)
where the subscript $(\cdot)_{phys}$ denotes physical quantities.
Let us hereafter use the shorter notation $N$ and $W_{f}$ for normalized corrected quantities given as
$N$ $=\frac{N_{corr}-O_{N}}{S_{N}},$ (23)
$W_{f}$ $=\frac{W_{f,corr}-O_{W}}{S_{W}},$ (24)
where $O_{N}$, $O_{W}$, $S_{N}$, and $S_{W}$ are the normalization offsets and scales for spool speed and fuel flow, respectively.
The simplified control-affine model of the GTE is thus given as
@@
where $f$ and $g$ are the autonomous and input dynamics, respectively.
@@
$J_{S}=$ $\left\|\dot{N}-\mathbf{\Xi_{f}}\mathbf{\Theta_{f}}(N,\mathbf{p_{f}})+\mathbf{\Xi_{g}}\mathbf{\Theta_{g}}(N,\mathbf{p_{g}})W_{f}\right\|_{F}^{2}$
$+\alpha\|\mathbf{\Xi}\|_{1},$ (26)
@@
Since also the nonlinear parameters of basis functions are optimized, this problem is solved using gradient descent (GD) methods with adaptive momentum (ADAM) optimizer [bib51]. Additionally, to promote sparsity of the solution, the L1 regularization is combined with sequential thresholding (ST) [bib24]. Whenever any weight in $\mathbf{\Xi}$ falls below a specified threshold, it is excluded from the equation. After each ST round, the GD algorithm is initialized again to fine-tune the remaining active terms in the model. This procedure repeats until a suitable trade-off between the complexity, given by the number of active terms, and accuracy is reached.
@@
$\begin{bmatrix}\mathbf{\Xi}\\ \textbf{p}\end{bmatrix}_{k+1}=\begin{bmatrix}\mathbf{\Xi}\\ \textbf{p}\end{bmatrix}_{k}-\left(\frac{\alpha\mathbf{v}}{\sqrt{\mathbf{s}+\epsilon}}\right)_{k},$ (27)
@@
$\textbf{v}_{k+1}=\beta_{1}\textbf{v}_{k}+(1-\beta_{1})\nabla J_{S,k},$ (28)
$\textbf{s}_{k+1}=\beta_{2}\textbf{s}_{k}+(1-\beta_{2}){\nabla J_{S,k}}^{2},$ (29)
$\alpha_{k}=\eta_{p}\left(\frac{1-{\beta_{2}}^{k}}{1-{\beta_{1}}^{k}}\right),$ (30)
$\nabla J_{S,k}=\begin{bmatrix}\frac{\partial J_{S}}{\partial\mathbf{\Xi}}&\frac{\partial J_{S}}{\partial\textbf{p}}\end{bmatrix}_{k}^{\intercal},$ (31)
@@
$\nabla J_{0}=K_{0}\frac{|f|}{f}\begin{bmatrix}\frac{\partial f}{\partial\mathbf{\Xi}}&\frac{\partial f}{\partial\textbf{p}}\end{bmatrix}^{\intercal}\bigg|_{t=0},$ (33)
@@
The ST is performed according to the following conditions:
$\xi=\begin{cases}0,\ \text{for }\ |\xi|\leq\xi_{min}\\ \xi,\ \text{for }\ |\xi|>\xi_{min}\end{cases},$ (34)
where $\xi_{min}$ is the threshold, and also according to the properties of the given basis functions. For instance, if logistic functions (LFs) were utilized, a shift of the center too far from the active region would make the function either 0 or 1, which would effectively exclude it or merge it with the other constant functions in the set.
Next step is the extraction of the autonomous dynamics $f(N)$ by setting $W_{f}=0\hskip 2.84526ptkg/s$, and generating a set of autonomous trajectories. The function $f(N)$ is integrated starting from multiple initial conditions until the zero state is reached. This provides an estimate of the true autonomous dynamics of the GTE. It should be noted that the zero normalized state, $N=0$, corresponds to the minimum sustainable physical spool speed, usually in the ground idle (GI) regime. The GI physical spool speed and fuel flow are subtracted from the corrected data, $N_{corr}$ and $W_{f,corr}$, and the result is scaled for the normalization according to (23) and (24). In this paper, the normalization was performed using $O_{N}=5000$ RPM, $S_{N}=10000$, $O_{W}=0.04$ kg/s, and $S_{W}=0.8$.